# Non-Euclidean Spatial Graph Neural Network

### 17 Dec 2023 v1

@@ 2 Related Work
Geometric deep learning. In the deep learning era, many research works tried to extend standard deep learning methods to geometric data such as manifolds and graphs [6]. Existing deep learning methods on manifolds can be categorized according to the way they describe 3D data formats [42, 36]. 3D ShapeNets[42] represents a geometric 3D shape as a probability distribution on a 3D voxel grid. Su et al. [36] construct multi-view representations of 3D shapes by using multiple images from different viewpoints. The main disadvantage of these methods is the efficiency of representing 3D data. Models are difficult to scale because they require much higher complexity. Alternatively, there exists another stream of works that represents manifold data by describing the 3D shape surface [29]. Pioneering works [31] generalize convolutional networks to non-Euclidean manifolds with a local geodesic system. Following works continue exploring the architecture of convolutions such as using localized spectral convolutional networks [44]. On the other hand, now there exist two significant streams of graph convolution. The first stream is spectral-based. Noticeable models such as ChebNet utilize spectral graph filters and their variants [37]. A main disadvantage of spectral approaches is that the spectral definition of convolution relies on the domain-dependent Fourier basis [6]. It is difficult to transfer a model learned on one graph to another graph with a different Fourier basis. The other stream of work is spatial-based, which directly performs convolution operations based in a message-passing manner. Methods such as GCN and GIN extract and aggregate neighborhood information for convolution [32, 43].
@@ 4 Method
(4.2) $\begin{split}d^{(k)}&=||\mathbf{l}^{(k)}||_{2},\theta^{(k)}=\arccos{\langle\frac{\mathbf{l}^{(k)}}{d^{(k)}},\frac{\mathbf{L}_{i,j}}{d_{i,j}}\rangle},\\ \quad\mathbf{L}_{i,j}&=\mathbf{p}_{j}-\mathbf{p}_{i},d_{i,j}=||\mathbf{L}_{i,j}||_{2}\\ \phi^{(k,k-1)}&=\langle\frac{\mathbf{c}^{(k-1)}\times\mathbf{c}^{(k)}}{||\mathbf{c}^{(k-1)}\times\mathbf{c}^{(k)}||_{2}},\frac{\mathbf{L}_{ij}}{||\mathbf{L}_{ij}||_{2}}\rangle\cdot\bar{\phi}^{(k,k-1)},\\ \phi^{(k,k+1)}&=\langle\frac{\mathbf{c}^{(k)}\times\mathbf{c}^{(k+1)}}{||\mathbf{c}^{(k)}\times\mathbf{c}^{(k+1)}||_{2}},\frac{\mathbf{L}_{ij}}{||\mathbf{L}_{ij}||_{2}}\rangle\cdot\bar{\phi}^{(k,k+1)},\\ \bar{\phi}^{(k,k-1)}&=\arccos\langle\mathbf{c}^{(k)},\mathbf{c}^{(k-1)}\rangle,\\ \bar{\phi}^{(k,k+1)}&=\arccos\langle\mathbf{c}^{(k)},\mathbf{c}^{(k+1)}\rangle,\\ \mathbf{c}^{(k)}=&\frac{\mathbf{L}_{ij}\times\mathbf{l}^{(k)}}{||\mathbf{L}_{ij}\times\mathbf{l}^{(k)}||_{2}},\varphi^{(k-1,k+1)}={\arccos{\langle\mathbf{n}^{(k-1)},\mathbf{n}^{(k+1)}\rangle}},\\ \mathbf{c}^{(k-1)}=&\frac{\mathbf{L}_{ij}\times\mathbf{l}^{(k-1)}}{||\mathbf{L}_{ij}\times\mathbf{l}^{(k-1)}||_{2}},\varphi^{(k-1,k)}={\arccos{\langle\mathbf{n}^{(k-1)},\mathbf{n}^{(k)}\rangle}},\\ \mathbf{c}^{(k+1)}=&\frac{\mathbf{L}_{ij}\times\mathbf{l}^{(k+1)}}{||\mathbf{L}_{ij}\times\mathbf{l}^{(k+1)}||_{2}},\varphi^{(k,k+1)}={\arccos{\langle\mathbf{n}^{(k)},\mathbf{n}^{(k+1)}\rangle}}\end{split}$
@@ 4.3 Complexity analysis
The time complexity of an $L$-layer GNN is $O(L\left|\mathcal{E}\right|b+L\left|\mathcal{V}\right|b^{2})$, where $b$ is the number of latent dimensions. Second, the time complexity of extracting geometric features from edge trajectory by LSTM is $O(L\left|\mathcal{E}\right|(Kb^{2}+Kbd))$ where $K$ is the average length of spatial trajectories and $d$ is the number of computed geometric features. Therefore, the total time complexity of our algorithm is $O(L\left|\mathcal{E}\right|(Kb^{2}+Kbd+b)+L\left|\mathcal{V}\right|b^{2})$.
@@ 5 Experimental Results
In this section, we first introduce the experimental settings, then the effectiveness of our proposed framework on both synthetic and real-world datasets is presented through a set of comprehensive experiments. The link to our code and appendix is at an anonymous GitHub repository https://github.com/anonymous-submitter-1234/manifold_spatial_networks/tree/main.
@@ 5.1.3 Comparison models
5.3 Roubustness Results.
@@
5.3.1 Effect of manifold irregularity analysis.

### 10 Jan 2024 v2

@@ 2 Related Work
Geometric deep learning. In the deep learning era, many research works tried to extend standard deep learning methods [21, 29, 20, 51, 9, 32, 45, 50] to geometric data such as manifolds and graphs [6]. Existing deep learning methods on manifolds can be categorized according to the way they describe 3D data formats [47, 40]. 3D ShapeNets[47] represents a geometric 3D shape as a probability distribution on a 3D voxel grid. Su et al. [40] construct multi-view representations of 3D shapes by using multiple images from different viewpoints. The main disadvantage of these methods is the efficiency of representing 3D data. Models are difficult to scale because they require much higher complexity. Alternatively, there exists another stream of works that represents manifold data by describing the 3D shape surface [33]. Pioneering works [35] generalize convolutional networks to non-Euclidean manifolds with a local geodesic system. Following works continue exploring the architecture of convolutions such as using localized spectral convolutional networks [49]. On the other hand, now there exist two significant streams of graph convolution. The first stream is spectral-based. Noticeable models such as ChebNet utilize spectral graph filters and their variants [41]. A main disadvantage of spectral approaches is that the spectral definition of convolution relies on the domain-dependent Fourier basis [6]. It is difficult to transfer a model learned on one graph to another graph with a different Fourier basis. The other stream of work is spatial-based, which directly performs convolution operations based in a message-passing manner. Methods such as GCN and GIN extract and aggregate neighborhood information for convolution [36, 48].
@@ 4 Method
$\begin{split}d^{(k)}&=||\mathbf{l}^{(k)}||_{2},\theta^{(k)}=\arccos{\langle\frac{\mathbf{l}^{(k)}}{d^{(k)}},\frac{\mathbf{L}_{i,j}}{d_{i,j}}\rangle},\\ \quad\mathbf{L}_{i,j}&=\mathbf{p}_{j}-\mathbf{p}_{i},d_{i,j}=||\mathbf{L}_{i,j}||_{2}\\ \phi^{(k,k-1)}&=\langle\frac{\mathbf{c}^{(k-1)}\times\mathbf{c}^{(k)}}{||\mathbf{c}^{(k-1)}\times\mathbf{c}^{(k)}||_{2}},\frac{\mathbf{L}_{ij}}{||\mathbf{L}_{ij}||_{2}}\rangle\cdot\bar{\phi}^{(k,k-1)},\\ \phi^{(k,k+1)}&=\langle\frac{\mathbf{c}^{(k)}\times\mathbf{c}^{(k+1)}}{||\mathbf{c}^{(k)}\times\mathbf{c}^{(k+1)}||_{2}},\frac{\mathbf{L}_{ij}}{||\mathbf{L}_{ij}||_{2}}\rangle\cdot\bar{\phi}^{(k,k+1)},\\ \bar{\phi}^{(k,k-1)}&=\arccos\langle\mathbf{c}^{(k)},\mathbf{c}^{(k-1)}\rangle,\\ \bar{\phi}^{(k,k+1)}&=\arccos\langle\mathbf{c}^{(k)},\mathbf{c}^{(k+1)}\rangle,\\ \mathbf{c}^{(k)}=&\frac{\mathbf{L}_{ij}\times\mathbf{l}^{(k)}}{||\mathbf{L}_{ij}\times\mathbf{l}^{(k)}||_{2}},\varphi^{(k-1,k+1)}={\arccos{\langle\mathbf{n}^{(k-1)},\mathbf{n}^{(k+1)}\rangle}},\\ \mathbf{c}^{(k-1)}=&\frac{\mathbf{L}_{ij}\times\mathbf{l}^{(k-1)}}{||\mathbf{L}_{ij}\times\mathbf{l}^{(k-1)}||_{2}},\varphi^{(k-1,k)}={\arccos{\langle\mathbf{n}^{(k-1)},\mathbf{n}^{(k)}\rangle}},\\ \mathbf{c}^{(k+1)}=&\frac{\mathbf{L}_{ij}\times\mathbf{l}^{(k+1)}}{||\mathbf{L}_{ij}\times\mathbf{l}^{(k+1)}||_{2}},\varphi^{(k,k+1)}={\arccos{\langle\mathbf{n}^{(k)},\mathbf{n}^{(k+1)}\rangle}}\end{split}$
@@ 4.3 Complexity analysis
The time complexity of an $L$-layer GNN is $O(L\left|\mathcal{E}\right|b+L\left|\mathcal{V}\right|b^{2})$, where $b$ is the number of latent dimensions. Second, the time complexity of extracting geometric features from edge trajectory by LSTM is $O(L\left|\mathcal{E}\right|(Kb^{2}+Kbd))$ where $K$ is the average length of spatial trajectories and $d$ is the number of computed geometric features. Therefore, the total time complexity of our algorithm is $O(L\left|\mathcal{E}\right|(Kb^{2}+Kbd+b)+L\left|\mathcal{V}\right|b^{2})$. The complete proof of Theorem A.1 is shown in the Appendix due to space limit. Theorem A.1 shows that the angle bound is defined by the plane angle and the circumradius.
@@ 5 Experimental Results
In this section, we first introduce the experimental settings, then the effectiveness of our proposed framework on both synthetic and real-world datasets is presented. The link to our code is at the GitHub repository https://github.com/rollingstonezz/SDM24_Manifold_spatial_networks.
@@ 5.2 Effectiveness Results
(4) It is also worth noting that on the SHREC dataset, only methods that consider orientation information achieved competitive results, which may arguably indicate that orientation information is important in the prediction task on this dataset.
@@
5.3 Effect of manifold irregularity analysis.